OpenStudy (anonymous):

The solutions of an absolute-value equation are given. What is the equation? x= b + or - 2a

OpenStudy (anonymous):

what does \(x=b+\) mean?

OpenStudy (anonymous):

ooh maybe it is \[x=b\pm2a\]

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

if you start with \[|x-b|=2a\] and solve for \(x\) you get \(x-b=2a\) or \(x-b=-2a\) so \(x=b\pm2a\)

OpenStudy (anonymous):

Wowza you're smart! Thank you, but how did you know to start with that equation?

OpenStudy (e.mccormick):

You gave \(x=b\pm 2a\) so s73 basically saw the answer. If you want to work for the answers, you do it in reverse. The \(\pm\) being the key point. All the \(\pm\) stuff must be on one side of the equation. Everything else on the other.

OpenStudy (anonymous):

Ohh gotcha that makes a lot of sense, thank you! I appreciate it!

OpenStudy (e.mccormick):

\(x=b\pm 2a\) OK, so just 2a is after the \(\pm\) It needs to be alone with its \(\pm\) \(x=b\pm 2a\implies \) \(x-b=\pm 2a\implies \) \(|x-b|= 2a \) This last step basically turning the \(\pm\) back into | |. So undoing math... which there is a lot of. =) Glad you got it!

OpenStudy (anonymous):

Thank you, I fanned ya!